Given a Solid 6061-T6 Member, place on torsion tester with pulley measuring degrees. Find the required weight to obtain a torque of 95 in*lb

Known Values:

Torque = 95 in*lb

Weight = "?"

Length of Member = 44.5 lb

Diameter = .5 inches

Radius of Pulley = 4.75 inches

Degrees of Twist from zero = 10 Degrees

Modulus of elasticity: 3.75 x 10^6 psi

Select from A)10 B)15 C) 20 D) 25
Answer #1

The front view of a torsion tester with pulley is shown below:

T r W

Pulley radius r = 4.75 inch

Required torque T = 95 in-lb

Weight = W

From the figure, T = W*r

therefore, W = T/r

W = 95/4.75 = 20lb (option C)